OpenStudy (anonymous):

Evaluate the integral: int (x^4 + 6x^3 + 10x^2 + x) dx / (x^2 + 6x +10) After dividing, I get int x^2 dx + int (x) dx / (x^2 + 6x +10) and this is where I am stuck...

OpenStudy (anonymous):

i don't get your answer.. \[\huge \int x^2 dx + \frac{\int xdx}{x^2 + 6x + 10}\] ????

OpenStudy (anonymous):

@xlegendx, \[\int\frac{x^4+6x^3+10x^2+x}{x^2+6x+10}~dx\] Finding the quotient via long division yields \[\int x^2~dx+\int\frac{x}{x^2+6x+10}~dx\]

OpenStudy (jhannybean):

Ahh, LD..

OpenStudy (anonymous):

yes sith... I don't understand how to take care of the second integral now...

OpenStudy (anonymous):

lol that makes more sense

OpenStudy (anonymous):

lol sorry

OpenStudy (anonymous):

if it were me, i would use algebraic substitution since the denominator seems to be prime

OpenStudy (jhannybean):

I think we'll have to use logarithms for the second part....just a hunch.

OpenStudy (anonymous):

substitution: so u = x^2 + 6x + 10 du = 2x + 6 dx but I have x not 2x + 6 right?

OpenStudy (anonymous):

then just subtract 6 from the numerator so nothing changes

OpenStudy (anonymous):

@gadav478, have you tried rewriting the integrand? \[\int\frac{x}{x^2+6x+10}~dx\] Note that letting \(u=x^2+6x+10\) gives you \(du=(2x+6)~dx\). So let's rewrite the integral: \[\frac{1}{2}\int\frac{2x+6-6}{x^2+6x+10}~dx\\ \frac{1}{2}\left[ \int\frac{2x+6}{x^2+6x+10}~dx - 6\int\frac{dx}{x^2+6x+10}~dx \right]\] Now applying that substitution: \[\frac{1}{2}\int\frac{du}{u} - 3\int\frac{dx}{x^2+6x+10}~dx\] For the second integral, complete the square in the denominator, then use a trig sub.

OpenStudy (anonymous):

see?

OpenStudy (anonymous):

ohhh

OpenStudy (jhannybean):

Where did this 1/2 outside the integrand come from....

OpenStudy (anonymous):

to neutralize 2x

OpenStudy (anonymous):

@Jhannybean, I multiplied the integral by 2/2, then distributed the numerator's 2.

OpenStudy (anonymous):

thanks so much everyone

OpenStudy (jhannybean):

Hmm... i'll go over this.

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

for anyone still here I have: x^3/3 + ln(x^2 +6x+10) - 6 arctan (x+3) + c

OpenStudy (anonymous):

forgot to distribute the 1/2: x^3/3 + ln(x^2 +6x+10) - 3 arctan (x+3) + c